Question 16 Tries remaining: 2 Points out of 6.25 Suppose there was an industry with 5 identical firms with marginal cost given by: MC- 10 + 2.5q. Use horizontal summation to find the total supply function for this industry. Fill in the blanks below. Flag question (Round to the nearest two decimal places if necessary. You will not need to include a negative sign in your response.,) Check
Answer #1

As we know MC is the supply curve of the firm and MC = P

Therefore, MC = P

P = 10 + 2.5Qi

2.5Qi = P - 10

Qi = 0.4P - 4

Individual firms supply is, Qi = 0.4P - 4

5 firms supply will be

Q = 5*(0.4P - 4)

Q = 2P - 20

Answer is 20.

Please contact if having any query thank you.